ganeshie8 (ganeshie8):

find the maximum value of \[25(a-b)^2 + (\sqrt{100-4a^2}+5\sqrt{1-b^2})^2\]

mathslover (mathslover):

Okay so I see some logic over here. The previous question that I asked was something similar to the expression, but we are required to find the maximum value here. Not a big deal. Let us try it out. Let \(\sqrt{100-4a^2} = y\) \(y^2 + 4a^2 = 100\) \(y^2 + (2a)^2 = 100\) Seems more like a circle with center \((y,2a)\) and radius \(10\) units. Now, \(5\sqrt{1-b^2} = x\) \(x^2 = 25(1 - b^2)\) \(x^2 = 25 - 25b^2\) \(x^2 + 25b^2 = 25\) \((x)^2 + (5b)^2 = 5^2\) o.O So we have got another equation for a circle, with center : \((x,5b)\) and radius \(5\) units.

ganeshie8 (ganeshie8):

thats an excellent approach @mathslover :) you're just 2 steps away from picking the right conics that work !

mathslover (mathslover):

So, I took right conics or not?

ganeshie8 (ganeshie8):

looks like you almost have them! but they still require some tweaking.. especially the 25 that is attached to the first term in the expression requires special care :P

mathslover (mathslover):

Hmm! :/ \((5a - 5b)^2 + (\sqrt{100-4a^2} + \sqrt{25(1-b^2)} )^2 \) Let \(y_1 = \sqrt{100 - 4a^2}\) \(\implies {y_1}^2 = 100 - 4a^2\) \({y_1}^2 + 4a^2 = 100 \) Let \(y_2 = \sqrt{5(1-b^2)}\) \(\implies {y_2}^2 = 25(1-b^2) \) \({y_2}^2 + 25b^2 = 25 \) Hmm! :/

mathslover (mathslover):

Am still trying... :(

mathslover (mathslover):

That "+" in the second term is annoying... \(25(a-b)^2 + (\sqrt{100-4a^2} \color{red}{+} 5\sqrt{1-b^2})^2\)

ganeshie8 (ganeshie8):

\[25\left[(a-b)^2 + \left(\dfrac{\sqrt{100-4a^2}}{5}+\sqrt{1-b^2}\right)^2\right]\]

ganeshie8 (ganeshie8):

spotting the conics might be simple now ?

OpenStudy (kainui):

My head is still spinning from this this morning. I'm gonna sit this out and watch.

mathslover (mathslover):

Cool! Let's see what's the difference: \((a-b)^2 + \left( \cfrac{\sqrt{100-4a^2} }{5} + \sqrt{1-b^2} \right)^2 \) I can take two obvious points from here, the x coordinates of both the points will be : \(a\) and \(b\) Let us now try to focus upon y -coordinates. Let \(\cfrac{\sqrt{100-4a^2}}{5} = y_1\) \(25{y_1}^2 = 100 - 4a^2\) \(25{y_1}^2 + 4a^2 = 100\) Cool! We have, a circle with center \((5 y_1 , 2a)\) and radius = 10 Next : Let \(y_2 = \sqrt{1-b^2} \\ {y_2}^2 = 1 - b^2 \\ {y_2}^2 + b^2 = 1 \) So, we have a circle with center \( (y_2 , b) \) and radius = 1 We've \(\left(a ,\cfrac{\sqrt{100-4a^2}}{5} \right) \) and \(\left(b , \sqrt{1-b^2} \right)\) So, we are trying to find maximum distance between two circles of given centers as \((5 y_1 , 2a)\) and \( (y_2 , b) \) respectively.

mathslover (mathslover):

Weird. I'm pretty sure that I'm doing something wrong. I will check everything again, one sec.

ganeshie8 (ganeshie8):

No ,you're doing great :) just this : \(25{y_1}^2 + 4a^2 = 100 \) is more like an ellipse... replacing \(a\) by \(x\) makes it easy to see : \[25y^2 + 4x^2 = 100 \] \[\dfrac{y^2}{4} + \dfrac{x^2}{25} = 1\]

mathslover (mathslover):

o.O :O O.o Didn't observe that ... nice..! So, we have one circle and one ellipse.. hmm!

ganeshie8 (ganeshie8):

the other conic is an unit circle as you guessed, but it is centered at origin !

mathslover (mathslover):

My concepts... :/ I need to go back and revise my concepts. :( Well, its really a nice question. I will try to proceed from here.

ganeshie8 (ganeshie8):

Okay... :) im not giving any useful hints as i don't want to spoil the fun for you !

OpenStudy (anonymous):

hmm then what just maximize x^2/25+y^2/4=1 with partial derivatives ?

OpenStudy (xapproachesinfinity):

how is y^2+4a^2=100 a circle?

ganeshie8 (ganeshie8):

it is an ellipse @xapproachesinfinity

ganeshie8 (ganeshie8):

25y^2+4a^2=100 y^2/4 + a^2/25 = 1

OpenStudy (xapproachesinfinity):

he got me lost when he said it is a circle lol

OpenStudy (xapproachesinfinity):

yeah i recognize the ellipse equation there

ganeshie8 (ganeshie8):

:) the question is still open, mathslover solved like 90% of it but its gona take a while for others to understand what he was doing...

OpenStudy (xapproachesinfinity):

i'm not really acquainted with ellipses and hyperbolas i need to review their equations ============= can we use calculus here?

ganeshie8 (ganeshie8):

any method is fine :) i feel calculus might give nasty derivatives for solving hmm

ganeshie8 (ganeshie8):

seeing the given expression as related to the `distance formula` is the key here, but im really looking forward to find ANY other cool ways to approach this problem xD

OpenStudy (anonymous):

i see then i'll work on it :)

mathslover (mathslover):

I got it.

mathslover (mathslover):

Is it 15? @ganeshie8

OpenStudy (anonymous):

ok here is one method ( reducing error to get high accurate ) |a|<=5 |b|<=1 and we need |a-b| br the largest difference , i'll stop here so other reveal more excited solution , looking forward for the geometrically solution sounds interesting !

mathslover (mathslover):

For the maximum distance ....

ganeshie8 (ganeshie8):

You have correct numbers for bounds @Marki ! xD

OpenStudy (anonymous):

;) also solution

ganeshie8 (ganeshie8):

one sec mathslover, i only saved the maximum value of given expression..

mathslover (mathslover):

Okay so here is what I've done: I've point as : \( \left(5a , \sqrt{100 -4a^2} \right) \) and \(\left(5b, -5\sqrt{1-b^2} \right) \) 2 Conics that we have are : \({y_1}^2 + 4a^2 = 100 \\ \cfrac{{y_1}^2}{100} + \cfrac{a^2}{25} = 1 \ \ \ \ \ \ \ \ \ [A]\) a = 10 and b = 5 And \(\cfrac{{y_2}^2}{25} + \cfrac{b^2}{1} = 1 \ \ \ [B]\) a = 5 and b = 1 So, we have two equations \([A]\) and \([B]\) for both ellipse. We need to find maximum distance between these ellipse which will be in fact sum of semi-major axis of larger ellipse (A) and semi-major axis of (B) So, we have maximum distance as : 10 + 5 = 15 units

mathslover (mathslover):



ganeshie8 (ganeshie8):

Excellent! but there are few numbers misplaced here and there - dont wry about it :) method is more important, not so much the final answer !

mathslover (mathslover):

I will check them too! :)

mathslover (mathslover):

One sec.

OpenStudy (kainui):

I've been waiting for the answer to this for a couple days now, I can't wait, this is exciting! =D

ganeshie8 (ganeshie8):

\[25\left[(a-b)^2 + \left(\dfrac{\sqrt{100-4a^2}}{5}+\sqrt{1-b^2}\right)^2\right]\] the stuff inside square brackets can be interpreted as the square of distance between below conics : \(y = \dfrac{\sqrt{100-4x^2}}{5}\) \(y = \sqrt{1-x^2}\) Look at the graph and eyeball the maximum value between them : http://gyazo.com/572956ac53a4dfb4a2884a451e28bf2b the maximum distance is 5--1 = 6 square it and multiply by 25 for the maximum value of the required expression

mathslover (mathslover):

Aw, can you help me in identifying the mistakes in my method?

ganeshie8 (ganeshie8):

there is one loophole in my solution though :P

ganeshie8 (ganeshie8):

@mathslover factoring out 25 will give you a circle and ellipse carrying out your exact same method yields the same answer given by wolfram :)

ganeshie8 (ganeshie8):

i don't see anything wrong in your steps..

mathslover (mathslover):

Yeah :/ As factoring out 25 will not make any differences in the final answer, isn't it? And from my answer, I've the maximum value as (15)^2 = 225... does this match to your final answer?

OpenStudy (kainui):

This seems like the hard way to do Lagrange multipliers lol.

mathslover (mathslover):

Cool! We just need to add 675 to my answer and you're done... LOL Not a big deal ... :P

OpenStudy (anonymous):

well as i got same to ganesh approch |a-b|=6 so 25(6^2)=900

ganeshie8 (ganeshie8):

Absolutely! not really a big deal at all haha ;) it would be more fun to find the minimum distance between a parabola and circle as eyeballing wont work in that case

OpenStudy (anonymous):

I totally agree =)